OpenStudy (anonymous):

John goes for a run. From his house, he jogs north for exactly 5.0 min at an average speed of 8.0 km/h. He continues north at a speed of 12.0 km/h for the next 30.0 min. He then turns around and jogs south at a speed of 15.0 km/h for 15.0 min. Then he jogs south for another 20.0 min at 8.0 km/h. He walks the rest of the way home.

OpenStudy (anonymous):

D = DS + DN = 20/3[km] + 77/12[km] = 157/12[km]. http://www.assignmentexpert.com/free-questions/question-on-physics-other-23367.html